# Dated Data: Tracing Knowledge Cutoffs in Large Language Models

### v1

@@ Abstract
Released Large Language Models (LLMs) are often paired with a claimed knowledge cutoff date, or the dates at which training data was gathered. Such information is crucial for applications where the LLM must provide up to date information. However, this statement only scratches the surface: do all resources in the training data share the same knowledge cutoff date? Does the model’s demonstrated knowledge for these subsets closely align to their cutoff dates? In this work, we define the notion of an effective cutoff. This is distinct from the LLM designer reported cutoff and applies separately to sub-resources and topics. We propose a simple approach to estimate effective cutoffs on the resource-level temporal alignment of an LLM by probing across versions of the data. Using this analysis, we find that effective cutoffs often differ from reported cutoffs. To understand the root cause of this observation, we conduct a direct large-scale analysis on open pre-training datasets. Our analysis reveals two reasons for these inconsistencies: (1) temporal biases of CommonCrawl data due to non-trivial amounts of old data in new dumps and (2) complications in LLM deduplication schemes involving semantic duplicates and lexical near-duplicates. Overall, our results show that knowledge cutoffs are not as simple as they have seemed and that care must be taken both by LLM dataset curators as well as practitioners who seek to use information from these models.
@@ 1 Introduction
Many Large Language Model (LLM) creators use the label “open-source” to describe their models, but very few actually provide the exact datasets their models used for pre-training. While the open-data community has made great efforts in training LLMs and providing the fully replicable training data (Groeneveld et al., 2024; Biderman et al., 2023; Computer, 2023), these models often lag behind their closed-data counterparts and are typically only used for research, leading most practitioners to use high-performing models whose data they cannot inspect.
In place of providing the exact pre-training data, creators often provide a “knowledge cutoff” date for their model. When faced with a description that states, e.g.,“this model has a cutoff date of March 2024,” does that mean all of its included resources share the exact same cutoff date? Even if the model provides an explicit reported cutoff for a resource (e.g. the Wikipedia dump date), does that imply that the model’s knowledge for that resource, or effective cutoff, is the same as the reported cutoff date? For LLM users, these questions can be crucial: imagine a layperson using an LLM for tax advice, without realizing that the effective cutoff of the tax code is 2022 and thus outdated – despite the fact that the reported cutoff is advertised as 2023 (Fig. 1).
As a result there has been a push for researchers to document their data (Mitchell et al., 2018; Pushkarna et al., 2022; Gebru et al., 2021; Luccioni et al., 2022), identify what data is in these models with membership inference tests (Carlini et al., 2022; Piktus et al., 2023; Marone & Van Durme, 2023), and otherwise reproduce data from their training set (Carlini et al., 2021; Ippolito et al., 2022; Nasr et al., 2023; Weller et al., 2023). However, these tests generally only check for static inclusion of the data, rather than identifying when the resources stopped being included. Even more difficult are cases where there exists multiple versions of a resource, where different versions can contain information that is updated, deleted, or even conflicting with the previous versions.
Given the crucial importance of these knowledge cutoffs and the lack of transparency from LLM creators, we seek to automatically determine the effective cutoff of models with respect to a given resource, without needing access to the model’s training data. Our contributions include: (1) we collect resource sets spanning long time frames and propose a simple method to determine effective cutoff of LLMs; (2) we show that for a variety of models (particularly newer models), the resource-level effective cutoffs differ drastically from their reported cutoff date; and (3) we provide analysis detailing the cause of such misalignment, showing that pre-training datasets suffer from deduplication complications and that CommonCrawl dumps exhibit temporal misalignment from the dump dates.
Figure 1: LLMs may contain different versions of a dataset in their training data than what is specified in a “cutoff" date, misleading users and causing errors when users are unaware of these differences.
@@ 2 Related Work
As the size of the data in LLMs increases, there have been many calls for researchers to document datasets through additional documentation artifacts. One of the earliest works on this topic calls for creators to broadly summarize the type of data trained on, calling these Model Cards (Mitchell et al., 2018). Other work proposed Data Cards and Datasheets that focus on more specific documentation of the data used (Pushkarna et al., 2022; Gebru et al., 2021). The open-access community has adopted versions of these (e.g. Huggingface model descriptions) but they do not provide fine-grained versioning and analysis that enables precise tracking of cutoffs. For example, it may not be clear whether scraped Common Crawl data contains additional versions of scraped Wikipedia.
Other research has focused on more fine-grained analysis, such as the role of filtering in LLM-data creation (Lucy et al., 2024) or how PII, toxic data, n-grams, and provenance play a role in LLM data (Dodge et al., 2021; Elazar et al., 2023). These works have provided great insight into LLMs, but necessarily depend on the data being available to the public. In contrast, our work focuses instead on determining what temporal versions of data exist in a model, without access to the training data.
Membership Inference Attacks
Many prominent LLMs (e.g. GPT-4, LLaMAv2) do not provide their training data or even descriptive information about them. In these cases, many people wish to know if their data is included in the model’s training set. Techniques like membership inference attacks (Shokri et al., 2017) have been applied to large language models. Many strategies have been proposed for this problem, most of which rely on the LLM’s perplexity over potential data instances: these include using similar but synthetic data (Mattern et al., 2023), prompting calibration (Fu et al., 2023), and a variety of other techniques (Shi et al., 2023; Faysse et al., 2024; Hisamoto et al., 2019).
Overall, membership inference testing focuses on whether a single instance was included in the LLMs training set (with the critical assumption that there was only one version). In contrast, our work focuses on which version(s) of the data was included in the LLMs training data.
@@
New written knowledge increases every day, but language models remain static. As re-training a LLM is prohibitively expensive, it is infeasible for LLMs to keep up with living online resources. [^1] For example, Wikipedia gets edited about every 2 seconds, with an average of 539 new articles a day. Thus, there exists a large field of research on continual learning, or helping LLMs stay up to date without expensive re-training. This typically involves modeling approaches that perform limited continued training to keep model knowledge up to date (e.g. Hu et al. 2023, among others; Kasai et al. 2023, among others). Our work also involves examining the temporal knowledge in these models, but differs by focusing on the original static models and what knowledge they contain.
@@ 3 Methodology
We seek to probe LLMs to determine their resource-level effective cutoffs. Models sometimes report an inclusion date for their sub-corpora,  [^2] https://huggingface.co/allenai/OLMo-7B reports using the January 2023 dump of Wikipedia which is typically the last timestamp of that resource but does not address additional sources of earlier data. This is relevant because given a certain Wikipedia dump is included in training data, it is reasonable to assume that the effective knowledge cutoff for those articles is the corresponding dump date. However, older versions of that text may be present in web scrapes in resources like the CommonCrawl.
To perform our analysis, we first gather several long spanning (2016 - 2023) resource collections. We then measure perplexity on data at certain dates across a variety of models and analyze the implications of these perplexity-at-time curves, referring back to pre-training data when available.
@@ 3.1 Time-Spanning Datasets
Broadly, online resources included in LLM training data can be divided into three subsets: resources that involve frequent updates (e.g. legal texts or Wikipedia), resources that are static but build over time (e.g. blogs or news articles), and purely static resources (e.g. books). We construct representative datasets that allow us to probe the former two types, choosing Wikipedia for our frequently updating resource and the New York Times for a building-over-time resource.
Wikipedia
Figure 2: Perplexity of the Wikipedia document “Liverpool” for Pythia-6.9B. Each point represents the perplexity of a version of the document “Liverpool” at the corresponding month.
Wikipedia is commonly used in pre-training (Gao et al., 2020; Almazrouei et al., 2023a), provides broad coverage over many topics, and changes frequently and rapidly. To create the dataset, we start by collecting the 5000333We filter out 100 that did not exist all the way back to 2016. most edited documents by number of edits. [^4] https://en.wikipedia.org/wiki/Wikipedia:Most_frequently_edited_pages, downloaded in May 2023 For each of these documents, we use the Wikipedia API to collect a version of the document at monthly intervals, from April 2016 to April 2023. Our final dataset therefore consists of the same 5000 documents changing monthly over this seven year period. We call this dataset WikiSpan. An individual document in WikiSpan is the Wikipedia document on topic $T$ at month $M$ (e.g. United States in January 2020).
New York Times
We use New York Times (NYT) articles to represent our building-over-time resource, as the documents contain long-form high quality text, are frequently included in pre-training data and CommonCrawl dumps, and provide a long and steady stream of new documents. For our probing dataset, called NewsSpan  we collect all the articles with top level domain “nytimes.com” from a curated collection of the 20 most recent CommonCrawl dumps (Soldaini et al., 2024). We bucket the articles by month according to their publish date, and collect 500 articles from each bucket from January 2016 until July 2020. Note that due to copyright concerns, CommonCrawl removed NYT articles from recent dumps and have stopped scraping it. Unlike the documents in WikiSpan, the documents in each bucket have no relation with one another.
@@
Pythia (Biderman et al., 2023) 2020
GPT-Neo (Black et al., 2022) 2020
GPT-J (Wang & Komatsuzaki, 2021) 2020
@@
Falcon (Almazrouei et al., 2023b) Mar ’20 Feb ’23
FalconRW (Almazrouei et al., 2023b) Feb ’23
@@
LLaMA (Touvron et al., 2023a) 5 (’17-’20) Aug ’22 2020
Table 1: Different decoder-only LLMs and their corresponding pre-training data. The CommonCrawl dumps are processed to various degrees. Frequently used datasets with their own columns include RefinedWeb (RW) (Penedo et al., 2023), C4 (Raffel et al., 2020), and the Pile (Gao et al., 2020).
@@ 3.2 Probing Methodology
We evaluate LLMs to automatically determine training data cutoffs. We do this by measuring the perplexity for the documents in each month bucket, using the first 512 tokens for Wikipedia and 256 for NYT (due to shorter documents). See Fig. 2 for an example on one Wikipedia topic.
@@
Our goal is to determine the effective cutoff of the model for a given resource, rather than focusing on individual topics or documents. Occasionally some documents in our time-spanning datasets have drastically different perplexity compared to previous months (e.g. a Wikipedia page changing to become a redirect page rather than a content page for one month). As this distracts from understanding the relative model perplexity across months, we follow previous work (Shi et al., 2023) and aggregate perplexities by taking the average of the median 95% of measurements in each month.
As perplexity measurements are not comparable across models (due to differences in tokenizers, etc.) we normalize the averaged perplexities to a 0-1 scale by performing min-max scaling in order to fairly compare their fluctuations across time. We call these relative perplexities.
@@ 3.3 Pre-Training Datasets
The LLMs we evaluate were pre-trained on data derived from three major datasets: C4 (Raffel et al., 2020), the Pile (Gao et al., 2020), and RefinedWeb (Penedo et al., 2023), as well as additional CommonCrawl dumps beyond those that were already included in these named datasets. We describe their contents, as well as how different LLMs used them during training. We note that no open pre-training dataset ever includes a direct dump of NYT articles; they are only present in included CommonCrawl dumps. While datasets may also contain other curated data (e.g. books), we focus on scraped data and collections that may contain Wikipedia or NYTimes content.
CommonCrawl
As the primary source of data for large language models, CommonCrawl (CC) dumps provide raw web text and metadata. Most recent models choose to use CCNet (Wenzek et al., 2019) which processes the files into higher quality text.
@@
C4 is a single, heavily processed, open-access CommonCrawl dump from April 2019. It took the processed data and applied content-based filters to discard documents containing undesirable content such as obscene words, boilerplate templates, and code. It is deduplicated at a three sentence span level, and has an overall size of about 750 GB.
@@
RefinedWeb (RW) consists of URLs from all CommonCrawl dumps at the time of their publication, from 2008 to February 2023. Because RW was designed to be used in conjunction with other high quality data sources, URLs from specific top-level domains (including “wikipedia.org") are excluded. The public RW is only a 600 billion token sample of the total 5 trillion token dataset.
@@ 3.4 Models
Pythia (Biderman et al., 2023), GPT-Neo (Black et al., 2022), and GPT-J (Wang & Komatsuzaki, 2021) are trained on The Pile (Gao et al., 2020). These represent some of the earliest attempts to replicate closed-access models by the open-source community. The Pythia suite was designed to provide a replicable training process for studying ideas like training dynamics, the impact of data, and memorization in LLMs. The RedPajama model suite (Computer, 2023) was intended as an open access replica of the LLaMA model (Touvron et al., 2023a), providing replicable pre-training data sourced from web-crawls. The Falcon and FalconRW suites (Almazrouei et al., 2023b) are trained on RefinedWeb, a dataset consisting solely of web-scraped data. Falcon also incorporates other curated corpora into its training data. Finally OLMo (Groeneveld et al., 2024), and LLaMA (Touvron et al., 2023a) are the latest generation of LLMs, trained on mixtures of the datasets used in previous systems and increasing amounts of CommonCrawl.
Overall, these models span several iterations of LLM pre-training approaches in terms of data selection, size of the training set, and model parameter size. Table 1 outlines these models suites and Section 3.3 describes corresponding pre-training datasets in detail.
3.5 Mining from Pre-training Data
We hypothesize that similar documents in training impact model knowledge and relative perplexity measurements. For example, if parts of a resource were duplicated many times, we might expect perplexity on those documents to be particularly low. Prior work has shown the effects of document frequency on LLM memorization (Carlini et al., 2022). To better understand the perplexity curves, we search for documents similar to those in our time-spanning datasets. This retrieves old versions of the documents, near duplicates, and copied fragments – all of which may impact information in the model and our perplexity measurements. We expect that the distribution mined from training data is inversely correlated with the perplexity trends – the counts of the versions of the retrieved Wikipedia-alike documents should be higher at the same months where perplexity is lower. The entire process consists of obtaining and indexing nearly 4T tokens from several LLM training sets.
Given a pre-training dataset, we first construct a BM25 index over it using Elasticsearch. For each topic, we use the first $k$ words of the document as the query to find similar documents. [^5] We use the version of Wikipedia from the model’s dump date for queries Note that BM25 score alone is not a reliable indicator of whether the document is a close duplicate (or even from Wikipedia) but is useful as a first step in filtering potential documents.
Using the top ten returned BM25 results, we calculate the edit distance between the matched documents and every version of the corresponding Wikipedia topic in WikiSpan, normalizing by the character length of the matched document to avoid length biases. We classify matched documents as a Wikipedia document only if the minimum of this normalized edit distance score is less than 0.2.666At most 20% of the document needs to be changed to exactly match a version of the corresponding topic.
With this subset of matched documents, we attribute each document to the month where its normalized edit distance was minimized (or multiple months if it is equally similar to many versions). This then allows us to plot the ground truth distribution of all documents similar to the original by date. We provide a more detailed description of this algorithm in Appendix A.

### v2

@@ Abstract
Large Language Models (LLMs) are often paired with a reported cutoff date, the time at which training data was gathered. Such information is crucial for applications where the LLM must provide up-to-date information. However, a reported cutoff only scratches the surface. Do all sub-resources in the training data share the same cutoff? Does the model’s demonstrated knowledge for these sub-resources closely align to their cutoff? We define the notion of an effective cutoff, which is distinct from the LLM’s reported cutoff and differs between sub-resources. We propose a simple approach to estimate effective cutoffs of an LLM on the resource-level by probing across versions of the data. Crucially, our method does not require access to a model’s pre-training data. Through our analysis, we find that effective cutoffs often drastically differ from reported cutoffs. To understand the root cause of this observation, we conduct a large-scale analysis on open pre-training datasets. Our analysis reveals two reasons for these inconsistencies: (1) temporal misalignments of CommonCrawl data due to non-trivial amounts of old data in new dumps; and (2) complications in LLM deduplication schemes involving semantic duplicates and lexical near-duplicates. Overall, our results show that cutoffs are not as simple as they have seemed and that care must be taken both by LLM dataset curators as well as practitioners who seek to use these models. We release our results and the code to replicate them at https://github.com/nexync/dated_data/.
@@ 1 Introduction
Many Large Language Model (LLM) creators do not elect to release their training data due to competitive reasons. In place of providing the exact pre-training data, they often provide a reported cutoff date for their model. When faced with a description that states, e.g.,“this model has a cutoff date of March 2024,” does that mean all of its included resources share the exact same cutoff date? Even if the model provides an explicit reported cutoff for a resource (e.g. the Wikipedia dump date), does that imply that the model’s knowledge of that resource, or effective cutoff, is the same as the reported cutoff date? For LLM users, these questions can be crucial: imagine a layperson using an LLM for tax advice, without realizing that the effective cutoff of the tax code is 2022 and thus outdated – despite the fact that the reported cutoff is advertised as 2023 (Fig. 1).
As a result, there has been a push for researchers to document their data (Mitchell et al., 2018; Pushkarna et al., 2022; Gebru et al., 2021; Luccioni et al., 2022), identify what data is in these models with membership inference tests (Carlini et al., 2022; Piktus et al., 2023; Marone & Van Durme, 2023), and otherwise reproduce data from their training set (Carlini et al., 2021; Ippolito et al., 2022; Nasr et al., 2023; Weller et al., 2023). However, these tests generally only check for static inclusion of the data, rather than identifying when the resources stopped being included. Even more difficult are cases where there exist multiple versions of a resource, where different versions can contain information that is updated, deleted, or even conflicting with the previous versions.
Given the crucial importance of these knowledge cutoffs and the lack of transparency from LLM creators, we seek to automatically determine the effective cutoff of models with respect to a given resource, without needing access to the model’s training data. We measure the perplexity of LLMs over varying versions of resources, identifying the effective cutoffs as the minima of the perplexity over time measurements. Our contributions are as follows: (1) we collect resource sets spanning long time frames and propose a simple method to determine effective cutoff of LLMs; (2) we show that for a variety of models (particularly newer models), the resource-level effective cutoffs differ drastically from their reported cutoff date; and (3) we provide an analysis detailing the causes of these misalignments, showing that pre-training datasets suffer from deduplication complications and that CommonCrawl dumps exhibit temporal misalignment from the dump dates.
Figure 1: LLMs may contain different versions of a dataset in their training data than what is specified in a “cutoff” date, misleading users and causing potential errors.
@@ 2 Related Work
As the size of the data in LLMs increases, there have been many calls for researchers to document datasets through additional documentation artifacts. These include Model Cards, Datasheets, and Data Cards – each focusing on documentation of a specific part of a model or data source (Mitchell et al., 2018; Gebru et al., 2021; Pushkarna et al., 2022). The open-access community has adopted versions of these (e.g. Huggingface model descriptions) but they do not provide fine-grained versioning that enables precise tracking of cutoffs. For example, it may not be clear whether scraped Common Crawl data contains additional versions of scraped Wikipedia.
Other research has focused on more fine-grained analysis, such as the role of filtering in LLM-data creation (Gururangan et al., 2022; Lucy et al., 2024) or how PII, toxic data, n-grams, and provenance play a role in LLM data (Dodge et al., 2021; Elazar et al., 2023). These works have provided great insight into LLMs, but necessarily depend on the data being available to the public. In contrast, our work focuses instead on determining what temporal versions of data exist in a model, without access to the training data.
Membership Inference
Many prominent LLMs do not provide their training data or even descriptive information about them, leading people to wonder if their data is included in the model’s training set. Techniques like membership inference attacks (Shokri et al., 2017) have been applied to LLMs. Many strategies have been proposed for this problem: they include using similar but synthetic data (Mattern et al., 2023), prompting calibration (Fu et al., 2023), and a variety of other techniques (Hisamoto et al., 2019; Shi et al., 2023; Faysse et al., 2024). While most strategies rely on the LLM’s perplexity over potential data instances, there has also been effort in black-box attacks using cloze tasks (Chang et al., 2023). Overall, membership inference testing focuses on whether an instance was included in the LLMs training set (with the critical assumption that there was only one version). In contrast, our work focuses on which version(s) of the data was included in the LLMs training data.
@@
New written knowledge increases every day, but language models remain static. As re-training a LLM is prohibitively expensive, it is infeasible for LLMs to keep up with living online resources. [^1] For example, Wikipedia gets edited about every 2 seconds Thus, there exists a large field of research on continual learning, or helping LLMs stay up to date without expensive re-training. This typically involves modeling approaches that perform limited continued training to keep model knowledge up to date (e.g. Hu et al. 2023, among others; Kasai et al. 2023, among others). Our work also involves examining the temporal knowledge in these models, but differs by focusing on if there is temporal misalignment in the original static models and what knowledge they contain rather than techniques to align them.
@@ 3 Methodology
We seek to probe LLMs to determine their resource-level effective cutoffs, defining the effective cutoff date of a model with respect to a resource as the date of the version of the resource that is most closely aligned with the model. This effective cutoff date can differ from the inclusion date of a model’s sub-resources that LLM creators sometime provide, which is typically the last timestamp of that resource but does not address additional sources of earlier data. [^2] https://huggingface.co/allenai/OLMo-7B reports using the January 2023 dump of Wikipedia This is relevant because given that a certain Wikipedia dump is included in training data, it is reasonable to assume that the effective cutoff for those articles is the corresponding dump date. However, older versions of similar text may be present in web scrapes like CommonCrawl.
To perform our analysis, we construct long spanning (2016-2023) datasets and measure perplexity on the data across a variety of models. We then analyze the implications of these perplexity-at-time curves and verify our results with the ground truth pre-training data.
@@ 3.1 Time-Spanning Datasets
Broadly, online resources included in LLM training data can be divided into three subsets: resources that involve frequent updates (e.g. legal texts or Wikipedia), resources that are static but build over time (e.g. blogs or news articles), and purely static resources (e.g. books). We construct representative datasets that allow us to probe the first two types, choosing Wikipedia for the updating resource and New York Times for the building resource.
WikiSpan
Figure 2: Perplexity of the Wiki document “Liverpool” under Pythia. Each point is the perplexity of the document at that time.
Wikipedia is commonly used in pre-training (Gao et al., 2020), provides broad topic coverage, and changes frequently. To create the dataset, we collect the 5000333We filter out 100 topics that did not exist all the way back to 2016. most edited documents by number of edits. [^4] https://en.wikipedia.org/wiki/Wikipedia:Most_frequently_edited_pages, from May 2023 For each of these, we use the Wiki API to collect a version of the document at monthly intervals, from April 2016 to April 2023. Our final dataset therefore consists of the same 5000 documents changing monthly over this seven year period. We call this dataset WikiSpan, and individual documents are the version of the Wikipedia document on topic $T$ at month $M$.
NewsSpan
We use New York Times (NYT) articles to represent our building-over-time resource, as the documents contain long-form high quality text, are frequently included in pre-training data and CommonCrawl dumps, and provide a long and steady stream of new documents. For our probing dataset, called NewsSpan, we collect all the articles with top level domain “nytimes.com” from a curated collection of the 20 most recent CommonCrawl dumps (Soldaini et al., 2024). We bucket the articles by month according to their publication date, and collect 500 articles from each bucket from January 2016 until July 2020. Note that due to copyright concerns, CommonCrawl removed NYT articles from recent dumps and have stopped scraping it. Unlike the documents in WikiSpan, the documents in each bucket have no relation with one another.
@@
Pythia (Biderman et al., 2023) ?  ’20
GPT-Neo (Black et al., 2022) ?  ’20
GPT-J (Wang & Komatsuzaki, 2021) ?  ’20
@@
Falcon (Almazrouei et al., 2023) Feb ’23
FalconRW (Almazrouei et al., 2023) Feb ’23
@@
LLaMA (Touvron et al., 2023a) 5 (’17-’20) Aug ’22 ?  ’20
Table 1: Different decoder-only LLMs and their corresponding pre-training data. The CommonCrawl dumps are processed to various degrees. Frequently used datasets with their own columns include RefinedWeb (RW), C4, and the Pile. CC Cutoff indicates the last CommonCrawl dump included, unknown months are marked with a ?.
@@ 3.2 Probing Methodology
Our goal is to determine the effective cutoff of the model for a given resource, rather than focusing on individual topics or documents. We do this by measuring perplexity on documents in each month bucket, using the first 512 tokens for Wikipedia and 256 for NYT (due to shorter documents). See Fig. 2 for an example.
@@
Occasionally some documents in our time-spanning datasets have drastically different perplexity compared to previous months (e.g. a Wikipedia page changing to become a redirect page rather than a content page for one month). As this distracts from understanding the relative model perplexity across months, we follow previous work and aggregate perplexities by taking an average of the measured perplexities after discarding the lowest and highest 2.5% of measurements in each month (Shi et al., 2023).
Perplexity measurements are not comparable across unrelated models due to data and training differences. Thus, we normalize the averaged perplexities to a 0-1 scale by performing min-max scaling over the entire time-span in order to compare their fluctuations across time. We call these relative perplexities. We take the time at which relative perplexities are minimized to be the effective knowledge cutoff. Since the minima are not always sharp, the effective cutoffs should be interpreted as a distribution over time.
3.3 Mining from Pre-training Data
We hypothesize that similar documents in training impact model knowledge and relative perplexity measurements. For example, if parts of a resource were duplicated many times, we might expect perplexity on those documents to be particularly low. Prior work has shown the effects of document frequency on LLM memorization (Carlini et al., 2022).
To better understand the perplexity curves, we search for documents similar to those in our time-spanning datasets. This retrieves old versions of the documents, near duplicates, and copied fragments – all of which may impact information in the model and our perplexity measurements. We expect that the distribution mined from training data is inversely correlated with the perplexity trends – the counts of the versions of the retrieved Wikipedia-alike documents should be higher at the same months where perplexity is lower. The entire process consists of obtaining and indexing nearly 4T tokens from several LLM training sets.
Given a pre-training dataset, we first construct a BM25 index over it using Elasticsearch. For each topic, we use the first 512 words of the document as the query to find similar documents. [^5] We use the version of Wikipedia from the model’s dump date for queries We use the BM25 scores as a first step filter in finding near duplicates.
Using the top ten BM25 results, we calculate the edit distance between the matched documents and every version of the corresponding Wikipedia topic in WikiSpan, normalizing by the character length of the matched document to avoid length biases. We classify matched documents as a Wikipedia document only if the minimum of this normalized edit distance score is less than 0.2.666At most 20% of a document can to be changed to count as an exact match a version of the topic. With this subset of similar documents, we attribute each document to its closest matching month by edit distance (including ties). This then allows us to plot the ground truth distribution of all documents similar to the original by date. We provide a more detailed description of this algorithm in Appendix A.
4 Experimental Setup
@@ 4.1 Pre-Training Datasets
The LLMs we evaluate were pre-trained on data derived from three major datasets: C4 (Raffel et al., 2020), the Pile (Gao et al., 2020), and RefinedWeb (Penedo et al., 2023), as well as additional CommonCrawl dumps. We describe their contents, as well as how different LLMs used them during training, focusing on subcorporas that may contain Wikipedia or NYTimes content. We note that no open pre-training dataset ever includes a direct dump of NYT articles; they are only present in included CommonCrawl dumps.
@@
C4 is a single, heavily processed, open-access CommonCrawl dump from April 2019. It uses content-based filters to discard documents containing undesirable content such as obscene words, boilerplate templates, and code. C4 is deduplicated at a three sentence span level, and has an overall size of about 750 GB.
@@
RefinedWeb (RW) consists of documents from CommonCrawl dumps spanning 2008 to February 2023. Because RW was designed to be used in conjunction with other high quality data sources, URLs from specific top-level domains (including “wikipedia.org”) are excluded. The public RW is only a 600B token sample of the total 5T token dataset.
@@ 4.2 Models
Pythia (Biderman et al., 2023), GPT-Neo (Black et al., 2022), and GPT-J (Wang & Komatsuzaki, 2021) were trained on the Pile, and represented early attempts to replicate closed-access models by the open-source community. The Pythia suite was designed to provide a replicable training process for studying ideas like training dynamics and memorization in LLMs. The RedPajama model suite (Computer, 2023) was intended as an open-access replica of the LLaMA model (Touvron et al., 2023a), providing replicable pre-training data sourced from web-crawls. The Falcon suites are trained on RefinedWeb, a dataset consisting solely of web-scraped data (Almazrouei et al., 2023). Finally, OLMo (Groeneveld et al., 2024), and LLaMA (Touvron et al., 2023a) are the latest generation of LLMs, trained on datasets used by prior models and large amounts of CommonCrawl. These models span several iterations of LLM pre-training approaches (data curation, data size, and model size). Table 1 describes these models in terms of their corresponding training sets and data versions.